OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (anonymous):

Are you really confused about how to do this, or are you just checking answers?

OpenStudy (anonymous):

i really have no idea how to do it. like honestly i don't.

OpenStudy (anonymous):

add 8 on the sides |2x – 3|=7 2x-3=7 and 2x-3=-7 now solve

OpenStudy (anonymous):

x=-2 or x=5